# Automated forest inventory: analysis of high-density airborne LiDAR point clouds with 3D deep learning

### 22 Dec 2023 v1

@@ 3.2 Tree features and stand attributes
For the prediction of tree crown attributes, i.e., crown diameter, crown volume, and crown volume of live branches, both accurate individual tree segmentation and individual tree component classification are required. For crown diameter prediction, RMIT and TUWIEN have relatively worse predictions 0.26 RMSEfor RMIT and 0.20 for TUWIEN, respectively. For example, two large, umbrella-shaped tree crowns in RMIT have been over-segmented, leading to incorrect crown dimaeter estimates, see Figure 7(a). Two examples from TUWIEN can be seen in Figure 7(b), where under-segmentation has led to large differences between estimated values and the reference. As the tree crowns in TUWIEN are large, inaccurately segmenting them causes significant mis-estimation of the crown diameter. Crown volumes (both over all branches and over only live branches) estimates are relatively poor for the RMIT and SCION regions. As shown in Figure 7 (examples 1 and 2), the estimates are again too low due to over-segmentation. For SCION (example 5) the predicted value is too large due to under-segmentation. In addition, tree component segmentation has a significant influence on the results: although the individual tree segmentation is good, some dead branches are mis-classified as live branches, leading to the final crown volume of live branches being greatly over-estimated. This exposes a conceptual issue of this common variable: the crown volume of a tree has a considerable definition uncertainty, and a minor change in what is regarded as the crown can lead to large changes of the volume.
@@ 4.1 Baseline comparisons
We have compared our proposed pipeline to three alternatives. Our first baseline is the best practice available in standard cloud processing software: Treeiso (Xi and Hopkinson, 2022) is a recently published, unsupervised algorithm to segment individual trees, which is available as a plug-in for the CloudCompare software (https://www.cloudcompare.org, last accessed 09/2023). Treeiso requires point clouds consisting exclusively of tree points. We therefore feed it only the points assigned to trees by our semantic segmentation module, under the TreeMix setting. The user parameters of Treeiso must be tuned for optimal performance, but are quite user-friendly and understandable also for non-experts in forestry.
@@
Table 7: Quantitative comparison to individual tree segmentation with Treeiso. All values are percentages (%). “default” are results obtained with the recommended default parameters of Treeiso, while “best” are those obtained with parameters that we individually tuned for each plot. Refer to Table 8 for those optimal values.
Method Completeness Omission error Commission error F-score
Treeiso (default) 64.1 35.9 43.1 60.3
Treeiso (best) 70.0 30.0 33.5 68.2
Ours 82.0 18.0 11.7 85.1
@@
Another comparison was made against a different deep learning-based method (Straker et al., 2023). That method involves constructing a CHM from the point cloud and converting it to pseudo-colour images, which are then fed into a 2D object detection network, YOLOv5 (Jocher et al., 2022). The quantitative comparison is given in Table 9. Our segmentation-based approach achieved higher F-scores in all forest regions except for CULS. Consistent with the findings reported by  Straker et al. (2023), both methods reach higher quality in coniferous forests, such as CULS, SCION, and NIBIO; whereas the quality somewhat deteriorates in forests with strongly varying tree heights and crown sizes, as in the RMIT and TUWIEN. Still, also in those challenging regions our segmentation method had an edge. On the particularly challenging TUWIEN plot that more than doubles the F-score from 30% to 69.4%.
Finally we also compare our individual tree segmentation results with to those of Point2Tree (Wielgosz et al., 2023), another graph-based method that was originally developed for terrestrial mobile laser scanning (MLS) point clouds. Point2Tree consists of two stages. First it employs a PointNet++ model to obtain semantic labels for the points. The semantic segmentation serves as a basis for an unsupervised, graph-based algorithm. That instance segmentation algorithm sequentially identifies individual trees in a graph: the segmented point cloud is divided into small segments that form the nodes of the graph, constructed based on trunk information. For optimal results, various hyper-parameters must be selected individually for each forest type, including slice size, minimum number of stem points, and stem height.
Point2Tree is particularly sensitive to low point density near the ground. This sensitivity arises because, having been developed for TLS points, graph construction starts from the bottom, and the PointNet++ has been trained on MLS data. If the point cloud is too sparse there, the method tends to break, which makes it less suitable for our ALS data. For instances with adequate density at lower height, Point2Tree is effective. When applied to For-Instance, the method only was successful for 3 of the test plots, shown in Table 10. We note that the failure cases could possibly be alleviated by re-training Point2Tree for ALS data and then adapting the graph based part to the new data characteristics.
@@ 4.3 Implications and limitations
The pipeline proposed in this paper sets a new state of the art for automated individual tree segmentation and semantic segmentation in ALS point clouds (at least for the forest characteristics covered by FOR-Instance). Based on the segmentation, it also delivers accurate estimates of key variables like tree height, crown width, and crown volume, as well as stand attributes like stem density and the terrain of the forest floor. These outputs provide a strong basis for a tree-level forest inventory, and consequently for fine-scale forest management. Complete per-tree information opens up the possibility to monitor tree growth dynamics over extended areas. It may also contribute to more targeted timber production, safeguarding of young trees, and maintenance of biodiversity within forest ecosystems. Moreover, the segmentation results may enhance the automatic estimation of ecosystem service indicators as well as aggregate timber volume, carbon stock and sequestration, in support of ecological management and protection.
@@
Of course, our proof-of-concept also has limitations. One important point is that it builds on fully supervised machine learning. Its performance is strongly heavily by the data seen during training, which means that it cannot necessarily be deployed out-of-the-box to unfamiliar forest types or sensor characteristics. Empirically, even moderate changes of the data characteristics can greatly affect neural network estimates. To sidestep extensive data collection efforts it may be useful to explore transfer learning and domain adaptation strategies (Triess et al., 2021). To that end, further empirical studies are needed, which in turn call for even larger and more diverse reference datasets.

### 23 Feb 2024 v2

@@ 3.2 Tree features and stand attributes
For the prediction of tree crown attributes, i.e., crown diameter, crown volume, and crown volume of live branches, both accurate individual tree segmentation and individual tree component classification are required. For crown diameter prediction, RMIT and TUWIEN have relatively worse predictions 0.26 RMSEfor RMIT and 0.20 for TUWIEN, respectively. For example, two large, umbrella-shaped tree crowns in RMIT have been over-segmented, leading to incorrect crown dimaeter estimates, see Figure 8(a). Two examples from TUWIEN can be seen in Figure 8(b), where under-segmentation has Δ=Δ Δ=Δ Δ=Δ Δ=Δ led to large differences between estimated values and the reference. As the tree crowns in TUWIEN are large, inaccurately segmenting them causes significant mis-estimation of the crown diameter. Crown volumes (both over all branches and over only live branches) estimates are relatively poor for the RMIT and SCION regions. As shown in Figure 8 (examples 1 and 2), the estimates are again too low due to over-segmentation. For SCION (example 5) the predicted value is too large due to under-segmentation. In addition, tree component segmentation has a significant influence on the results: although the individual tree segmentation is good, some dead branches are mis-classified as live branches, leading to the final crown volume of live branches being greatly over-estimated. This exposes a conceptual issue of this common variable: the crown volume of a tree has a considerable definition uncertainty, and a minor change in what is regarded as the crown can lead to large changes of the volume.
@@ 4.1 Baseline comparisons
We have compared our proposed pipeline to three alternatives. Our first baseline Δ=Δ Δ=Δ Δ=Δ Δ=Δ is an example of current standard practice, available in popular cloud processing software: Treeiso (Xi and Hopkinson, 2022) is a recently published, unsupervised algorithm to segment individual trees, which is available as a plug-in for the CloudCompare software (https://www.cloudcompare.org, last accessed 09/2023). Treeiso requires point clouds consisting exclusively of tree points. We therefore feed it only the points assigned to trees by our semantic segmentation module, under the TreeMix setting. The user parameters of Treeiso must be tuned for optimal performance, but are quite user-friendly and understandable also for non-experts in forestry.
@@
Table 8: Δ=Δ Δ=Δ Δ=Δ Δ=Δ Quantitative comparison to individual tree segmentation with Treeiso. All values are percentages (%). “default” are results obtained with the recommended default parameters of Treeiso, while “best” are those obtained with parameters that we individually tuned for each plot. Refer to Table 9 for those optimal values.
Method Completeness Omission error Commission error F-score Cov
Treeiso (default) 64.1 35.9 43.1 60.3 67.8
Treeiso (best) 70.0 30.0 33.5 68.2 71.7
Ours 82.0 18.0 11.7 85.1 78.1
@@
Another comparison was made against a different deep learning-based method (Straker et al., 2023). That method involves constructing a CHM from the point cloud and converting it to pseudo-colour images Δ=Δ Δ=Δ Δ=Δ Δ=Δ of 640 $\times$ 640 pixels, which are then fed into a 2D object detection network, YOLOv5 (Jocher et al., 2022). Δ=Δ Δ=Δ Δ=Δ Δ=Δ They used the published weights of the YOLOv5l-seg model and chose the same hyperparameters as described by Jocher et al. (2022), except for changing the initial learning rate to 0.001. The quantitative comparison is given in Table 10. Our segmentation-based approach achieved higher F-scores in all forest regions except for CULS. Consistent with the findings reported by  Straker et al. (2023), both methods reach higher quality in coniferous forests, such as CULS, SCION, and NIBIO; whereas the quality somewhat deteriorates in forests with strongly varying tree heights and crown sizes, as in the RMIT and TUWIEN. Still, also in those challenging regions our segmentation method had an edge. On the particularly challenging TUWIEN plot that more than doubles the F-score from 30% to 69.4%.
Finally we also compare our individual tree segmentation results Δ=Δ Δ=Δ Δ=Δ Δ=Δ to those of Point2Tree (Wielgosz et al., 2023), another graph-based Δ=Δ Δ=Δ Δ=Δ Δ=Δ method that was originally developed for terrestrial mobile laser scanning (MLS) point clouds. Point2Tree consists of two stages. First it employs a PointNet++ model to obtain semantic labels for the points. The semantic segmentation serves as a basis for an unsupervised, graph-based algorithm. That instance segmentation algorithm sequentially identifies individual trees in a graph: the segmented point cloud is divided into small segments that form the nodes of the graph, constructed based on trunk information. For optimal results, various hyper-parameters must be selected individually for each forest type, including slice size, minimum number of stem points, and stem height.
Point2Tree is particularly sensitive to low point density near the ground. This sensitivity arises because, having been developed for TLS points, graph construction starts from the bottom, and the PointNet++ has been trained on MLS data. If the point cloud is too sparse there, the method tends to break, which makes it less suitable for our ALS data. For instances with adequate density at lower height, Point2Tree is effective. When applied to FOR-Instance, the method was only successful for 3 of the test plots, shown in Table 11 Δ=Δ Δ=Δ Δ=Δ Δ=Δ ; showing that our method, adapted to the characteristics of ALS data, indeed offers a significant advantage when it comes to handling plots with varying point densities. For completeness we note that the failure cases could possibly be alleviated by re-training Point2Tree for ALS data and then adapting the graph -based part to the new data characteristics.
@@
Our method treats dominant and understory trees in a unified manner, in principle ensuring equal effectiveness for both. Still, understory trees may exhibit lower point density and suffer from missing data caused by severe occlusions, due to the nature of airborne sensor data acquisition. Consequently, their segmentation is usually less accurate compared to canopy-level trees (Dong et al., 2020; Jarron et al., 2020; Wang et al., 2023b). We are unable to quantify the difference in segmentation performance between dominant and understory trees for our pipeline, because of the lack of precise ground truth annotations in the dataset. This uncertainty again emphasizes the need for more complete and accurate ground truth. Reference datasets with carefully annotated understory trees will be a indispensable to overcome the current limitations and unlock the full potential of segmentation algorithms for forestry.
@@ 4.3 Implications and limitations
The pipeline proposed in this paper sets a new state of the art for automated individual tree segmentation and semantic segmentation in ALS point clouds (at least for the forest characteristics covered by FOR-Instance Δ=Δ Δ=Δ Δ=Δ Δ=Δ , including boreal coniferous forest, temperate coniferous forest, temperate deciduous forest and sclerophyll forest).
Δ=Δ Δ=Δ Δ=Δ Δ=Δ Our method has also demonstrated potential for detecting understory trees. While contributing minimally to biomass, they play a crucial role for the ecosystem by ensuring canopy succession, stand development, soil erosion protection, and habitat for wildlife (Hamraz et al., 2017). The detection of understory trees has recently received more attention within the research community (Hamraz et al., 2017; Dong et al., 2020; Jarron et al., 2020; Wang et al., 2023b; Penner et al., 2023). Current methods often follow a stratified approach that iteratively removes higher canopy layers, then detects lower ones (Dong et al., 2020; Jarron et al., 2020; Wang et al., 2023b). In contrast, our method directly processes 3D point clouds, without the need to derive a 2.5D height model and to update it for each layer. We argue that directly analyzing the 3D forest structure is a more natural approach to effectively and efficiently detect understory trees. Moreover, advances in LiDAR technology (e.g., single-photon LiDAR) can be expected to further increase point cloud density and to promote the adoption of 3D methods.
@@
Based on the segmentation, it also delivers accurate estimates of key variables like tree height, crown width, and crown volume, as well as stand attributes like stem density and the terrain of the forest floor. These outputs provide a strong basis for a tree-level forest inventory, and consequently for fine-scale forest management. Complete per-tree information opens up the possibility to monitor tree growth dynamics over extended areas. It may also contribute to more targeted timber production, safeguarding of young trees, and maintenance of biodiversity within forest ecosystems. Moreover, the segmentation results may enhance the automatic estimation of ecosystem service indicators as well as aggregate timber volume, carbon stock and sequestration, in support of ecological management and protection.
Of course, our proof-of-concept also has limitations. Δ=Δ Δ=Δ Δ=Δ Δ=Δ In particular, it tends to reach lower segmentation accuracy for more complex forest structure. Among the forest types contained in the FOR-Instance dataset, the CULS region consists solely of even-aged trees of a single tree species (Scots pine). The NIBIO area also features mature forest, dominated mainly by Picea abies. Although the SCION region has high tree density ($\approx$930 trees per hectare), it is a productive plantation forest also dominated by a single species (Pinus radiata). The segmentation accuracy in these three regions is high, which can be attributed to their relatively uniform species composition, with even-aged and mature trees creating a regular, single-layered forest structure. Despite the overlapping and intertwined crowns of the closely spaced, radiata pines with many branches, the segmentation accuracy in the SCION area remains high.
On the contrary, the TUWIEN and RMIT regions are structurally more complex. TUWIEN features complicated vegetation structures (woody debris, lying and standing deadwood, and dense understory), a large diversity of tree species, and considerable variations in canopy layering and tree size (DBH up to 1.3 m). RMIT encompasses trees with a wide range of ages and heights (from 5 m to 17 m), with an understory consisting of low shrubs ($\leq$2 m in height) and native grasses. For these two structurally more intricate forests our method exhibits markedly lower accuracy (Figure 6). It is worth noting again that TUWIEN and RMIT both only have a single training plot in the current training data. It is thus not clear to what extent the lower performance stems from the higher complexity and not simply from insufficient training data. It will be interesting to see how 3D deep learning copes with more varied forests once sufficiently large training sets become available.
Another factor is point cloud density. We have conducted a simple experiment to study the impact of point density on the segmentation accuracy of ForAINet: The original FOR-Instance data is sub-sampled to 7 different point densities (i.e., 10, 25, 50, 75, 100, 500, and 1000 pts/m2). That range of densities encompasses the typical densities of our ALS-HD (500-1000 pts/m2), through conventional, airborne high-density ALS (up to 100 pts/m2), to traditional ALS (approximately 10 pts/m2). We then run our best configuration (with TreeMix augmentation, see Section 3.1) on the sub-sampled datasets. The results indicate that the performance decreases with point density (Figure 10). The drop in instance segmentation F-score remains within 5 percentage points for densities $\geq$75 pts/m2. Below that value the omission error markedly increases. The mIoU for semantic segmentation follows a similar pattern. In conclusion, our 3D learning method is challenged by point densities below 100 pts/m2, as obtained for instance from traditional airborne ALS.
@@
Δ=Δ Δ=Δ Δ=Δ Δ=Δ Another important point is that Δ=Δ Δ=Δ Δ=Δ Δ=Δ our method builds on fully supervised machine learning. Its performance is strongly heavily by the data seen during training, which means that it cannot necessarily be deployed out-of-the-box to unfamiliar forest types Δ=Δ Δ=Δ Δ=Δ Δ=Δ (e.g., tropical environments with high overlap between trees) or sensor characteristics. Empirically, even moderate changes of the data characteristics can greatly affect neural network estimates. To sidestep extensive data collection efforts it may be useful to explore transfer learning and domain adaptation strategies (Triess et al., 2021). To that end, further empirical studies are needed, which in turn call for even larger and more diverse reference datasets.
Figure 10: Δ=Δ Δ=Δ Δ=Δ Δ=Δ Variations in segmentation metrics across point densities.